Cycloid can be represented by parametric equation.So what is its perimeter?

Grade:Upto college level

3 Answers

abhishek sahu,IIT-K
13 years ago

for more info--plz see-http://en.wikipedia.org/wiki/Cycloid

rajender kumar malik
13 years ago

cycloid is the curve defined by the path of a point on the edge of circular wheel as the wheel rolls along a straight line. It is an example of a roulette, a curve g
enerated by a curve rolling on another curve.

x=r(t-sint)

y=r(1-cost)

wher t is a parameter

                          DO APPROVE THE ANSWER IF IT HELPED YOU

 

 

 

 

 

 

 

 

 

 

 

Siddhartha Goswami
3 years ago
cycloid is the curve defined by the path of a point on the edge of circular wheel as the wheel rolls along a straight line
Its equation in polar form is 
=r(1-cosx)x=r(1-sinx)y